# Cohomology and Deformations of left-symmetric Rinehart Algebras

### v3

@@ 6.3 Trivial deformation
We study deformations of left-symmetric Rinehart algebra using the deformation cohomology. Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra, and $\mathfrak{m}\in C^{2}(L,L)$. Consider a $t$-parameterized family of multiplications $\mathfrak{m}_{t}:L[[t]]\otimes L[[t]]\rightarrow L[[t]]$ and linear maps $\ell_{t}:L\to Der(A)$ given by
$\mathfrak{m}_{t}(x,y)=$ $x\cdot y+t\mathfrak{m}(x,y),$ (6.9)
$\ell_{t}=$ $\ell+t\Xi_{\mathfrak{m}}$ (6.10)
@@
which implies that conditions (3.1) and (3.2) in Definition 3.1 are satisfied for $\mathfrak{m}_{t}$. Then we can deduce that $(L,A,\mathfrak{m}_{t},\ell_{t})$ is a deformation of $(L,A,\cdot,\ell)$ if and only if
@@
$=\mathfrak{m}(y,x\cdot z)-\mathfrak{m}(x,y\cdot z)-\mathfrak{m}([x,y],z),$ (6.11)
@@
$\mathfrak{m}(\mathfrak{m}(x,y),z)-\mathfrak{m}(x,\mathfrak{m}(y,z))$ $=$ $\mathfrak{m}(\mathfrak{m}(y,x),z)-\mathfrak{m}(y,\mathfrak{m}(x,z)).$ (6.12)
@@
$\mathfrak{m}(x,y)$ $=$ $x\cdot N(y)+N(x)\cdot y-N(x\cdot y),$ (6.13)
$N\mathfrak{m}(x,y)$ $=$ $N(x)\cdot N(y),,$ (6.14)
$\ell\circ N$ $=$ $\Xi_{\mathfrak{m}},.$ (6.15)
@@
$N(x)\cdot N(y)-x\cdot N(y)-N(x)\cdot y+N^{2}(x\cdot y)=0.$ (6.16)
@@
We have seen that a trivial deformation of a left-symmetric Rinehart algebra give rise to a Nijenhuis operator. In fact, the converse is also true as can be seen from the following theorem.
@@
Let $N$ be a Nijenhuis operator on a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$. Then a deformation of $(L,A,\cdot,\ell)$ can be obtained by putting
$\mathfrak{m}(x,y)=\delta(N)(x,y).$
@@
Proof. Since $\mathfrak{m}$ is exact, $\mathfrak{m}$ is closed naturally, i.e. Eq. (6.3) holds. To see that $\mathfrak{m}$ generates a deformation, we only need to show that (6.12) holds, which follows from the Nijenhuis condition (6.16). We omit the details. At last, it is obvious that
$(\mathrm{Id}+tN)(x\cdot_{t}y)=(\mathrm{Id}+tN)(x)\cdot(\mathrm{Id}+tN)(y),\quad\ell\circ(\mathrm{Id}+tN)=a_{t},$
@@
Let $N$ be a Nijenhuis operator on a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$. Then $(L,A,\cdot_{N},\ell_{N}=\ell\circ N)$ is a left-symmetric Rinehart algebra, where
@@
Proof. It is obvious to show that $(L,\cdot_{N})$ is a left-symmetric algebra and $\ell_{N}$ is a representation of L on $Der(A)$. Evidently, we have
$\ell_{N}(ax)=a\ell_{N}(x),\forall x\in L,a\in A..$
@@
Let $N$ be a Nijenhuis operator on a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$. Then for arbitrary positive $j,k\in\mathbb{N}$, the following equation holds
$N^{j}(x)\cdot N^{k}(y)-N^{k}(N^{j}(x)\cdot y)-N^{j}(x\cdot N^{k}(y))+N^{j+k}(x\cdot y)=0,\quad\forall~{}x,y\in L.$ (6.17)
@@
Let $(L,A,\cdot,\ell)$ left-symmetric Rinehart algebra and $N$ a Nijenhuis operator.
@@
Let $N$ be a Nijenhuis operator on a left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$. Then the operator $P(N)=\sum_{i=0}^{n}c_{i}N^{i}$ is a Nijenhuis operator. If $N$ is invertible, then $Q(N)=\sum_{i=-m}^{n}c_{i}N^{i}$ is also a Nijenhuis operator.
Proof. According to Lemma 6.16, we obtain
@@
$=$ $\sum_{i,j=0}^{n}c_{j}c_{k}\Big{(}N^{j}(x)\cdot N^{k}(y)-N^{k}(N^{j}(x)\cdot y)-N^{j}(x\cdot N^{k}(y))+N^{j+k}(x\cdot y)\Big{)}=0,\ \ \ \forall x,y\in L,$
@@ 7 $\mathcal{O}$-operators and Nijenhuis operators
In this section, we highlight the relationships between $\mathcal{O}$-operators and Nijenhuis operators on left-symmetric Rinehart algebra. Moreover, we illustrate some conections between Nijenhuis operators and compatible $\mathcal{O}$-operators on left-symmetric Rinehart algebra.
@@ 7.1 Relationships between $\mathcal{O}$-operators and Nijenhuis operators
We first give the definitions of an $\mathcal{O}$-operator and of Rota-Baxter opeartor.
@@
$T(au)$ $=$ $aT(u),$ (7.1)
$T(u)\cdot T(v)$ $=$ $T\Big{(}\rho(T(u))(v)+\mu(T(v))(u)\Big{)},\quad\forall u,v\in M,a\in A.$ (7.2)
@@
$\mathcal{R}(ax)$ $=$ $a\mathcal{R}(x),$ (7.3)
$\mathcal{R}(x)\cdot\mathcal{R}(y)$ $=$ $\mathcal{R}(\mathcal{R}(x)\cdot y+x\cdot\mathcal{R}(y)),\quad\forall x,y\in L,a\in A,$ (7.4)
@@
Proof. For Item (i), for all $x,y\in L$ then we have
$(N-Id)(x)\cdot(N-Id)(y)-(N-Id)\big{(}(N-Id)(x)\cdot y+x\cdot(N-Id)(y)\big{)}+2(N-Id)(x\cdot y)$
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $(M;\rho,\mu)$ be a representation on $L$. Let $T:M\rightarrow L$ be a linear map. For any $\lambda$, $T$ is an $\mathcal{O}$-operator on $L$ associated to $(M;\rho,\mu)$ if and only if the linear map $\mathcal{R}_{T,\lambda}:=\left(\begin{array}[]{cc}0&T\\ 0&-\lambda{\rm Id}\end{array}\right)$ is a Rota-Baxter operator of weight $\lambda$ on the semidirect product left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$, where the multiplication $\cdot_{L\oplus M}$ is given by (4.2).
Proof. It is easy to check the Eq.(7.3). Let $x_{1},x_{2}\in L$ and $m_{1},m_{2}\in M$,
@@
$=T(m_{1})\cdot T(m_{2})-\lambda\rho(T(m_{1})m_{2}-\lambda\mu(T(m_{2}))m_{1}.$ (7.5)
@@
$\lambda\mathcal{R}_{T,\lambda}((x_{1}+m_{1})\cdot_{L\oplus M}(x_{2}+m_{2}))$
@@
$\lambda\mathcal{R}_{T,\lambda}(x_{1}\cdot x_{2}+\rho(x_{1})m_{2}+\mu(x_{2})m_{1})$
$=$ $\mathcal{R}_{T,\lambda}\big{(}T(m_{1})\cdot x_{2}+\rho(T(m_{1}))m_{2}-\lambda\mu(x_{2})m_{1}+x_{1}\cdot T(m_{2})-\lambda\rho(x_{1})m_{2}+\mu(T(m_{2}))m_{1}\big{)}$
@@
$=$ $T\big{(}\rho(T(m_{1}))m_{2}+\mu(T(m_{2}))m_{1}\big{)}-\lambda T\big{(}\mu(x_{2})m_{1}+\rho(x_{1})m_{2}\big{)}-\lambda\Big{(}\rho(T(m_{1}))m_{2}+\mu(T(m_{2}))m_{1}\Big{)}$
$+\lambda^{2}\Big{(}\mu(x_{2})m_{1}+\rho(x_{1})m_{2}\big{)}+\lambda\big{(}T(\rho(x_{1})m_{2})+T(\mu(x_{2})m_{1})\big{)}-\lambda^{2}\big{(}\rho(x_{1})m_{2}+\mu(x_{2})m_{1})\big{)}$
$=$ $T\big{(}\rho(T(m_{1}))m_{2}+\mu(T(m_{2}))m_{1}\big{)}-\lambda\rho(T(m_{1}))m_{2}-\lambda\mu(T(m_{2}))m_{1}.$ (7.6)
According to Eqs. (7.5) and (7.6), $\mathcal{R}_{T,\lambda}$ is a Rota-Baxter operator of weight $\lambda$ on the semidirect product left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$ if and only if $T$ is an $\mathcal{O}$-operator on $(L,A,\cdot,\ell)$ associated to $(M;\rho,\mu)$.
@@
    $\mathcal{N}_{T}:=\left(\begin{array}[]{cc}0&T\\ 0&{\rm Id}\end{array}\right)$ is a Nijenhuis operator on left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$.
@@
Proof. Note that $\mathcal{N}_{T}=\mathcal{R}_{T,-1}$ and $(\mathcal{N}_{T})^{2}=\mathcal{N}_{T}$ then $\mathcal{N}_{T}$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$ using (iii) in Proposition 7.3.
Similarly $N_{T}=\mathcal{R}_{T,0}$ and $(N_{T})^{2}=0$, then $N_{T}$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$ according (ii) in Proposition 7.3.
@@ 7.2 Compatible $\mathcal{O}$-operators and Nijenhuis operators
Two operators $T_{1}$ and $T_{2}$ are compatible if and only if the equation (7.7) holds.
Proof. For all $u,v\in M$, $a\in A$, we have
@@
$-T_{1}\big{(}\rho(T_{1}(u))(v)+\mu(T_{1}(v))(u)\big{)}-T_{1}\big{(}\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u)\big{)}$
$-T_{2}\big{(}\rho(T_{1}(u))(v)+\mu(T_{1}(v))(u)\big{)}-T_{2}\big{(}\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u)\big{)}$
$=$ $T_{1}(u)\cdot T_{2}(v)+T_{2}(u)\cdot T_{1}(v)-T_{1}\Big{(}\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u)\Big{)}-T_{2}\Big{(}\rho(T_{1}(u))(v)+\mu(T_{1}(v))(u)\Big{)}$
Then $T_{1}+T_{2}$ is an $\mathcal{O}$-operator associated to $(M;\rho,\mu)$ if and only if Eq.(7.7) holds.
@@
The equation (7.7) implies that for any $k_{1},k_{2}$ the linear combination $k_{1}T_{1}+k_{2}T_{2}$ is an $\mathcal{O}$-operator.
@@
Proof. For all $x,y\in L$, since $T_{2}$ is invertible, there exist $u,v\in M$ such that $T_{2}(u)=x,T_{2}(v)=y$. Hence $N=T_{1}\circ T_{2}^{-1}$ is a Nijenhuis operator if and only if the following equation holds:
$NT_{2}(u)\cdot NT_{2}(v)=N(NT_{2}(u)\cdot T_{2}(v)+T_{2}(u)\cdot NT_{2}(v))-N^{2}(T_{2}(u)\cdot T_{2}(v)).$ (7.8)
Since $T_{1}=N\circ T_{2}$ is an $\mathcal{O}$-operator, the left hand side of the above equation is
@@
Hence Eq.(7.8) holds by acting $N$ on both sides of the last equality.
@@
The following results are an immediate consequence of the last two propositions.
@@
Since a Rota-Baxter operator of weight zero is an $\mathcal{O}$-operator on a left-symmetric Rinheart algebra associated to the adjoint representation, we have the following corollary.
Corollary 7.12.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinheart algebra. Suppose that $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are two invertible Rota-Baxter operators of weight zero. Then $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are compatible in the sense that any linear combination of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ is still a Rota-Baxter operator of weight zero if and only if $N=\mathcal{R}_{1}\circ\mathcal{R}_{2}^{-1}$ is a Nijenhuis operator.

### v5

@@ 6.3 Trivial deformation
We study deformations of left-symmetric Rinehart algebras using the deformation cohomology. Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra, and $\mathfrak{m}\in C^{2}(L,L)$. Consider a $t$-parameterized family of multiplications $\mathfrak{m}_{t}:L[[t]]\otimes L[[t]]\rightarrow L[[t]]$ and linear maps $\ell_{t}:L\to Der(A)$ given by
$\mathfrak{m}_{t}(x,y)=$ $x\cdot y+t\mathfrak{m}(x,y),$ (24)
$\ell_{t}=$ $\ell+t\Xi_{\mathfrak{m}}.$ (25)
@@
which implies that conditions (4) and (5) in Definition 3.1 are satisfied for $\mathfrak{m}_{t}$. Then we can deduce that $(L,A,\mathfrak{m}_{t},\ell_{t})$ is a deformation of $(L,A,\cdot,\ell)$ if and only if
@@
$=\mathfrak{m}(y,x\cdot z)-\mathfrak{m}(x,y\cdot z)-\mathfrak{m}([x,y],z),$ (26)
@@
$\mathfrak{m}(\mathfrak{m}(x,y),z)-\mathfrak{m}(x,\mathfrak{m}(y,z))$ $=$ $\mathfrak{m}(\mathfrak{m}(y,x),z)-\mathfrak{m}(y,\mathfrak{m}(x,z)).$ (27)
@@
$\mathfrak{m}(x,y)$ $=$ $x\cdot N(y)+N(x)\cdot y-N(x\cdot y),$ (28)
$N\mathfrak{m}(x,y)$ $=$ $N(x)\cdot N(y),$ (29)
$\ell\circ N$ $=$ $\Xi_{\mathfrak{m}}.$ (30)
@@
$N(x)\cdot N(y)-x\cdot N(y)-N(x)\cdot y+N^{2}(x\cdot y)=0.$ (31)
@@
We have seen that a trivial deformation of a left-symmetric Rinehart algebra gives rise to a Nijenhuis operator. In fact, the converse is also true as can be seen from the following theorem.
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N$ be a Nijenhuis operator. Then a deformation of $(L,A,\cdot,\ell)$ can be obtained by putting
$\mathfrak{m}(x,y)=\delta N(x,y).$
@@
Proof 6.18.
Since $\mathfrak{m}$ is a coboundary, then it is a cocycle, i.e. equation (6.3) holds. To see that $\mathfrak{m}$ generates a deformation, we only need to show that (27) holds, which follows from the Nijenhuis condition (31). At the end, we can easily check that
$(\mathrm{Id}+tN)(x\cdot_{t}y)=(\mathrm{Id}+tN)(x)\cdot(\mathrm{Id}+tN)(y),\quad\ell\circ(\mathrm{Id}+tN)=\ell_{t},$
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N$ be a Nijenhuis operator. Then $(L,A,\cdot_{N},\ell_{N}=\ell\circ N)$ is a left-symmetric Rinehart algebra, where
@@
Proof 6.20.
It is obvious to show that $(L,\cdot_{N})$ is a left-symmetric algebra and $\ell_{N}$ is a representation of L on $Der(A)$. Evidently, we have
$\ell_{N}(ax)=a\ell_{N}(x),\forall x\in L,a\in A.$
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N$ be a Nijenhuis operator. Then for arbitrary positive $j,k\in\mathbb{N}$, the following equation holds
$N^{j}(x)\cdot N^{k}(y)-N^{k}(N^{j}(x)\cdot y)-N^{j}(x\cdot N^{k}(y))+N^{j+k}(x\cdot y)=0,\quad\forall~{}x,y\in L.$ (32)
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N$ a Nijenhuis operator.
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $N$ be a Nijenhuis operator. Then the operator $P(N)=\sum_{i=0}^{n}c_{i}N^{i}$ is a Nijenhuis operator. If $N$ is invertible, then $Q(N)=\sum_{i=-m}^{n}c_{i}N^{i}$ is also a Nijenhuis operator.
Proof 6.24.
According to Lemma 6.21, we obtain, $\forall x,y\in L,$
@@
$=$ $\sum_{i,j=0}^{n}c_{j}c_{k}\Big{(}N^{j}(x)\cdot N^{k}(y)-N^{k}(N^{j}(x)\cdot y)-N^{j}(x\cdot N^{k}(y))+N^{j+k}(x\cdot y)\Big{)}=0,$
@@ 7 $\mathcal{O}$-operators and Nijenhuis operators
In this section, we highlight the relationships between $\mathcal{O}$-operators and Nijenhuis operators on left-symmetric Rinehart algebras. Moreover, we illustrate some connections between Nijenhuis operators and compatible $\mathcal{O}$-operators on left-symmetric Rinehart algebras.
@@ 7.1 Relationships between $\mathcal{O}$-operators and Nijenhuis operators
We first give the definitions of an $\mathcal{O}$-operator and of Rota-Baxter operator.
@@
$T(au)$ $=$ $aT(u),$ (33)
$T(u)\cdot T(v)$ $=$ $T\Big{(}\rho(T(u))(v)+\mu(T(v))(u)\Big{)},\quad\forall u,v\in M,a\in A.$ (34)
@@
$\mathcal{R}(ax)$ $=$ $a\mathcal{R}(x),$ (35)
$\mathcal{R}(x)\cdot\mathcal{R}(y)$ $=$ $\mathcal{R}(\mathcal{R}(x)\cdot y+x\cdot\mathcal{R}(y)),\quad\forall x,y\in L,a\in A,$ (36)
@@
Proof 7.4.
For Item (i), for all $x,y\in L$ then we have
$(N-Id)$ $(x)\cdot(N-Id)(y)$
$\qquad-(N-Id)\big{(}(N-Id)(x)\cdot y+x\cdot(N-Id)(y)\big{)}+2(N-Id)(x\cdot y)$
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $(M;\rho,\mu)$ be a representation on $L$. Let $T:M\rightarrow L$ be a linear map. For any $\lambda$, $T$ is an $\mathcal{O}$-operator on $L$ associated to $(M;\rho,\mu)$ if and only if the linear map $\mathcal{R}_{T,\lambda}:=\left(\begin{array}[]{cc}0&T\\ 0&-\lambda{\rm Id}\end{array}\right)$ is a Rota-Baxter operator of weight $\lambda$ on the semidirect product left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$, where the multiplication $\cdot_{L\oplus M}$ is given by (9).
Proof 7.6.
It is easy to check the equation (35). Let $x_{1},x_{2}\in L$ and $m_{1},m_{2}\in M$,
@@
$=T(m_{1})\cdot T(m_{2})-\lambda\rho(T(m_{1})m_{2}-\lambda\mu(T(m_{2}))m_{1}.$ (37)
@@
$\quad\lambda\mathcal{R}_{T,\lambda}((x_{1}+m_{1})\cdot_{L\oplus M}(x_{2}+m_{2}))$
@@
$\quad\lambda\mathcal{R}_{T,\lambda}(x_{1}\cdot x_{2}+\rho(x_{1})m_{2}+\mu(x_{2})m_{1})$
$=$ $\mathcal{R}_{T,\lambda}\big{(}T(m_{1})\cdot x_{2}+\rho(T(m_{1}))m_{2}-\lambda\mu(x_{2})m_{1}$ (38)
$\quad\qquad+x_{1}\cdot T(m_{2})-\lambda\rho(x_{1})m_{2}+\mu(T(m_{2}))m_{1}\big{)}$
@@
$=$ $T\big{(}\rho(T(m_{1}))m_{2}+\mu(T(m_{2}))m_{1}\big{)}-\lambda T\big{(}\mu(x_{2})m_{1}+\rho(x_{1})m_{2}\big{)}$ (39)
$-\lambda\Big{(}\rho(T(m_{1}))m_{2}+\mu(T(m_{2}))m_{1}\Big{)}+\lambda^{2}\Big{(}\mu(x_{2})m_{1}+\rho(x_{1})m_{2}\big{)}$
$+\lambda\big{(}T(\rho(x_{1})m_{2})+T(\mu(x_{2})m_{1})\big{)}-\lambda^{2}\big{(}\rho(x_{1})m_{2}+\mu(x_{2})m_{1})\big{)}$
$=$ $T\big{(}\rho(T(m_{1}))m_{2}+\mu(T(m_{2}))m_{1}\big{)}-\lambda\rho(T(m_{1}))m_{2}-\lambda\mu(T(m_{2}))m_{1}.$ (40)
According to equations (37) and (40), $\mathcal{R}_{T,\lambda}$ is a Rota-Baxter operator of weight $\lambda$ on the semidirect product left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$ if and only if $T$ is an $\mathcal{O}$-operator on $(L,A,\cdot,\ell)$ associated to $(M;\rho,\mu)$.
@@
    $\mathcal{N}_{T}:=\left(\begin{array}[]{cc}0&T\\ 0&{\rm Id}\end{array}\right)$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$.
@@
Proof 7.8.
Note that $\mathcal{N}_{T}=\mathcal{R}_{T,-1}$ and $(\mathcal{N}_{T})^{2}=\mathcal{N}_{T}$, thus $\mathcal{N}_{T}$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$, using (iii) in Proposition 7.3.
Similarly $N_{T}=\mathcal{R}_{T,0}$ and $(N_{T})^{2}=0$, then $N_{T}$ is a Nijenhuis operator on the left-symmetric Rinehart algebra $({L\oplus M},\cdot_{L\oplus M})$, according to (ii) in Proposition 7.3.
@@ 7.2 Compatible $\mathcal{O}$-operators and Nijenhuis operators
Two operators $T_{1}$ and $T_{2}$ are compatible if and only if the equation (41) holds.
Proof 7.11.
For all $u,v\in M$, $a\in A$, we have
@@
$\qquad-T_{1}\big{(}\rho(T_{1}(u))(v)+\mu(T_{1}(v))(u)\big{)}-T_{1}\big{(}\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u)\big{)}$
$\qquad-T_{2}\big{(}\rho(T_{1}(u))(v)+\mu(T_{1}(v))(u)\big{)}-T_{2}\big{(}\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u)\big{)}$
$=T_{1}(u)\cdot T_{2}(v)+T_{2}(u)\cdot T_{1}(v)$
$\qquad-T_{1}\Big{(}\rho(T_{2}(u))(v)+\mu(T_{2}(v))(u)\Big{)}-T_{2}\Big{(}\rho(T_{1}(u))(v)+\mu(T_{1}(v))(u)\Big{)}$
Then $T_{1}+T_{2}$ is an $\mathcal{O}$-operator associated to $(M;\rho,\mu)$ if and only if equation (41) holds.
@@
Equation (41) implies that for any $k_{1},k_{2}$ the linear combination $k_{1}T_{1}+k_{2}T_{2}$ is an $\mathcal{O}$-operator.
@@
Proof 7.14.
For all $x,y\in L$, since $T_{2}$ is invertible, there exist $u,v\in M$ such that $T_{2}(u)=x,$ $T_{2}(v)=y$. Hence $N=T_{1}\circ T_{2}^{-1}$ is a Nijenhuis operator if and only if the following equation holds:
$NT_{2}(u)\cdot NT_{2}(v)=N(NT_{2}(u)\cdot T_{2}(v)+T_{2}(u)\cdot NT_{2}(v))-N^{2}(T_{2}(u)\cdot T_{2}(v)).$ (42)
Since $T_{1}=N\circ T_{2}$ is an $\mathcal{O}$-operator, the left-hand side of the above equation is
@@
Hence, equation (42) holds by acting $N$ on both sides of the last equality.
@@
The following result is an immediate consequence of the last two propositions.
@@
Acknowledgement
The authors are thankful to the Deanship of Graduate Studies and Scientific Research at University of Bisha for supporting this work through the Fast-Track Research Support Program. Mohamed Elhamdadi was partially supported by Simons Foundation collaboration grant 712462.